I'm trying to help my friends kid with math homework and its been a while since I've done anything with graphs. The problem is...
The organizers of a walkathon are getting cost estimates from 2 printing companies for brochures to advertise the event.
C=cost in $ and n=#of brochures
Company A C=15+0.10n
Company B C=.25n
a. For what number of brochures are the costs the same for both companies, and what method did you use?
c. What info does the y-intercept represent for each equation?
I remembered enough to help her with b and d, but I'm lost on these two problems. I just can't remember how to find the answer. Thank you for your time.
Clark

Answer by Alan3354(69443)

C=cost in $ and n=#of brochures
Company A C=15+0.10n
Company B C=.25n
a. For what number of brochures are the costs the same for both companies, and what method did you use?
C = 15 + 0.10n = 0.25n when they cost is equal
15 + 0.10n = 0.25n
15 = 0.15n
n = 100
The total cost for 100 is the same for either price schedule.
-----------------------
c. What info does the y-intercept represent for each equation?
Assuming y is the cost, and x is the number of brochures,
The y-intercept is the cost of zero (0) brochures, or the base cost.
--------------